OpenStudy (anonymous):

sin(75°)cos(15°)+cos(75°)sin(15°) Simplify

OpenStudy (anonymous):

this is the addition angle formula for sine, so it is the same as \[\sin(75+15)=\sin(90)\]

OpenStudy (anonymous):

thank you!